OpenStudy (anonymous):

I give MEADLS! 70 - 4 × (7 - 5) to the power of 4 18 82 0 6

OpenStudy (anonymous):

U have to use BODMAS...[ Bracket; Of; division; multiplication; addition; subtraction] This is the order in which u have to do it..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the to the power four what is it for?

OpenStudy (anonymous):

is i t for (7-5) ?

OpenStudy (anonymous):

it goes to the top hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

\[(7-5) ^{4} ?\]

OpenStudy (anonymous):

yeah yeah

OpenStudy (anonymous):

wat yeah yeah? is it inside the bracket or outsidE?

OpenStudy (anonymous):

outside

OpenStudy (anonymous):

ok so ..7-5=2 2^4=16 ok?

OpenStudy (anonymous):

(7-5)^4=16 4*16=64 70-64=6

OpenStudy (anonymous):

then 4 X 16=64 70-64=6 understood?

OpenStudy (anonymous):

yeah thank you should you get the MEADL or him @Nikhil619

OpenStudy (anonymous):

give him

OpenStudy (anonymous):

IDK..it depend on you...and i dont really care..

OpenStudy (anonymous):

ok @Nikhil619 it is